# When to Trust: A Causality-Aware Calibration Framework for Accurate Knowledge Graph Retrieval-Augmented Generation

### v1

@@ Abstract.
$c$s: Information systems Language modelsccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Causal reasoning and diagnostics
@@ 1. Introduction
Figure 1. Calibration comparison on the MetaQA dataset. Calibration error quantifies the average discrepancy between a model’s predicted confidence and its actual correctness. The naive KG-RAG framework (a) is consistently over-confident with a high calibration error, whereas our Ca2KG framework (b) achieves much better calibration with a reduced error.
@@
Although KG-RAG has been shown to substantially improve factual accuracy in complex tasks (Xiang et al., 2025; Kim et al., 2023; Liu et al., 2024; Xia et al., 2025), the naive KG-RAG framework often exhibits such over-confidence despite factual inaccuracies. (see Figure 1). This issue directly reflects poor calibration, which measures how well a model’s predicted confidence aligns with its actual likelihood of being correct. A well-calibrated model outputs high confidence only when its predictions are reliable, and low confidence when uncertainty is warranted. Calibration is especially critical for KG-RAG because inaccurate confidence can mislead downstream reasoning steps, amplify retrieval errors, and cause incorrect entity linking or graph traversal. Therefore, improving calibration is urgently needed for reliable web-scale knowledge retrieval and generation.
@@ 3.4. Panel-based Re-scoring
You are given 3 candidate answers for the same query from different prompts: $t_{0}$: $a_{0}$; $t_{1}$: $a_{1}$; $t_{2}$: $a_{2}$. Task: (1) Build a unified set of UNIQUE answers by merging semantically identical strings across $a_{0}$, $a_{1}$, $a_{2}$. (2) Compute a GLOBAL frequency vector over the unique answers based on how many of $a_{0}$$a_{1}$$a_{2}$ map to each canonical answer. (3) ASSIGN probabilities for EACH interventions ($t_{0}$, $t_{1}$, $t_{2}$) to be EXACTLY this global frequency vector normalised by 3 (counts/3), after duplicate aggregation and before rounding. Re-scoring Structure: Merging Rules Ignore case, whitespace, punctuation, trivial formatting, and plural/singular differences. Choose a clean canonical form for answers (e.g., title case). If multiple inputs ($a_{0}$$a_{1}$$a_{2}$) map to the same canonical answer, aggregate them by SUMMING that answer’s frequency before normalisation. Probability rules (HARD CONSTRAINTS) Let count[i] be how many of $a_{0}$$a_{1}$$a_{2}$ map to answers[i]. Then for every intervention T in $t_{0}$, $t_{1}$, $t_{2}$, set T[i] = round(count[i]/3, 2). Do NOT output equal probabilities across answers when counts differ. Each probability $\in[0.00,1.00]$; sums may be ¡ 1.00 after rounding. If an intervention has no plausible answers, use zeros. Return EXACTLY one line of STRICT JSON, no wrapper, no extra text.
@@ Appendix B Case Study
Case Study I on WebQSP Question: what is the capital of argentina? Ground-truth: Buenos Aires Step 1: Counterfactual Prompting initial prompt $t_{0}$:
@@
Step 2: Panel-based Re-scoring You are given 3 candidate answers for the same question from different prompts: $t_{0}$: $a_{0}$; $t_{1}$: $a_{1}$; $t_{2}$: $a_{2}$. Task: (1) Build a unified set of UNIQUE answers by merging semantically identical strings across $a_{0}$, $a_{1}$, $a_{2}$. (2) Compute a GLOBAL frequency vector over the unique answers based on how many of $a_{0}$$a_{1}$$a_{2}$ map to each canonical answer. (3) Assign probabilities for each scenario ($t_{0}$, $t_{1}$, $t_{2}$) to be exactly this global frequency vector normalised by 3 (counts/3), after duplicate aggregation and before rounding. Merging rules: ignore case, whitespace, punctuation, and plural/singular differences; choose a clean canonical form for answers; if multiple inputs map to the same canonical answer, aggregate by summing before normalisation. Probability rules: A) Let count[i] be how many of $a_{0}$, $a_{1}$, $a_{2}$} map to answers[i]. Then for every scenario S in $t_{0}$, $t_{1}$, $t_{2}$}, set S[i] = round(count[i]/3, 2). B) Do not output equal probabilities across answers when counts differ. C) Each probability $\in[0.00,1.00]$; sums may be <1.00<1.00 after rounding. D) If a scenario has no plausible answers, use zeros. Return exactly one line of strict JSON, no wrapper, no extra text. Canonical merge: Buenos Aires, Argentina
@@
"Buenos Aires": 0.62, "Argentina": 0.38 Step 3: Accurate Answer Generation Panel JSON:
@@
Causal effect variation ($\text{CE}_{\text{var}}$): $\text{CE}_{\text{var}}$ $(\text{Buenos Aires})$ $=\max(0.62,0.68,0.80)-\min(0.62,0.68,0.80)$ $=0.80-0.62=0.18$ $\text{CE}_{\text{var}}$ $(\text{Argentina})$ $=\max(0.38,0.32,0.20)-\min(0.38,0.32,0.20)$ $=0.38-0.20=0.18$ Average causal effect ($\overline{\text{CE}}$): $\overline{\text{CE}}(\text{Buenos Aires})=(0.62+0.68+0.80)/3=0.70$ $\overline{\text{CE}}(\text{Argentina})=(0.38+0.32+0.20)/3=0.30$ Causal calibration index (CCI): $\text{CCI}(\text{Buenos Aires})=0.70\times(1-0.18)=0.70\times 0.82=0.57$ $\text{CCI}(\text{Argentina})=0.30\times(1-0.18)=0.30\times 0.82=0.25$ Final Answer: {Buenos Aires}
Case Study II on MetaQA Question: The films that share directors with the films [Tiresia] are written by who? Ground-truth: Bertrand Bonello Step 1: Counterfactual Prompting initial prompt $t_{0}$:
@@
Step 2: Panel-based Re-scoring You are given 3 candidate answers for the same question from different prompts: $t_{0}$: $a_{0}$; $t_{1}$: $a_{1}$; $t_{2}$: $a_{2}$. Task: (1) Build a unified set of UNIQUE answers by merging semantically identical strings across $a_{0}$, $a_{1}$, $a_{2}$. (2) Compute a GLOBAL frequency vector over the unique answers based on how many of $a_{0}$$a_{1}$$a_{2}$ map to each canonical answer. (3) Assign probabilities for each scenario ($t_{0}$, $t_{1}$, $t_{2}$) to be exactly this global frequency vector normalised by 3 (counts/3), after duplicate aggregation and before rounding. Merging rules: ignore case, whitespace, punctuation, and plural/singular differences; choose a clean canonical form for answers; if multiple inputs map to the same canonical answer, aggregate by summing before normalisation. Probability rules: A) Let count[i] be how many of $a_{0}$, $a_{1}$, $a_{2}$} map to answers[i]. Then for every scenario S in $t_{0}$, $t_{1}$, $t_{2}$}, set S[i] = round(count[i]/3, 2). B) Do not output equal probabilities across answers when counts differ. C) Each probability $\in[0.00,1.00]$; sums may be <1.00<1.00 after rounding. D) If a scenario has no plausible answers, use zeros. Return exactly one line of strict JSON, no wrapper, no extra text. Canonical merge: Bertrand Bonello, Luca Fazzi
@@
"Bertrand Bonello": 0.67, "Luca Fazzi": 0.33 Step 3: Accurate Answer Generation Panel JSON:
@@
Causal effect variation ($\text{CE}_{\text{var}}$): $\text{CE}_{\text{var}}$ $(\text{Bertrand Bonello})$ $=\max(0.67,0.67,0.67)-\min(0.67,0.67,0.67)$ $=0.00$ $\text{CE}_{\text{var}}$ $(\text{Luca Fazzi})$ $=\max(0.33,0.33,0.33)-\min(0.33,0.33,0.33)$ $=0.00$ Average causal effect ($\overline{\text{CE}}$): $\overline{\text{CE}}(\text{Bertrand Bonello})$ $=(0.67+0.67+0.67)/3=0.67$ $\overline{\text{CE}}(\text{Luca Fazzi})$ $=(0.33+0.33+0.33)/3=0.33$ Causal calibration index (CCI): $\text{CCI}(\text{Bertrand Bonello})$ $=0.67\times(1-0.00)=0.67$ $\text{CCI}(\text{Luca Fazzi})$ $=0.33\times(1-0.00)=0.33$ Final Answer: {Bertrand Bonello}

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792358isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Language modelsccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Causal reasoning and diagnostics
@@ 1. Introduction
Figure 1. Calibration comparison on the MetaQA dataset. Calibration error quantifies the average discrepancy between a model’s predicted confidence and its actual correctness. The naïve KG-RAG framework (a) is consistently over-confident with a high calibration error, whereas our Ca2KG framework (b) achieves much better calibration with a reduced error.
@@
Although KG-RAG has been shown to substantially improve factual accuracy in complex tasks (Xiang et al., 2025; Kim et al., 2023; Liu et al., 2024; Xia et al., 2026), the naïve KG-RAG framework often exhibits such over-confidence despite factual inaccuracies (see Figure 1). This issue directly reflects poor calibration, which measures how well a model’s predicted confidence aligns with its actual likelihood of being correct. A well-calibrated model outputs high confidence only when its predictions are reliable, and low confidence when uncertainty is warranted. Calibration is especially critical for KG-RAG because inaccurate confidence can mislead downstream reasoning steps, amplify retrieval errors, and cause incorrect entity linking or graph traversal. Therefore, improving calibration is urgently needed for reliable web-scale knowledge retrieval and generation.
@@ 3.4. Panel-based Re-scoring
You are given 3 candidate answers for the same query from different prompts: $t_{0}$: $a_{0}$; $t_{1}$: $a_{1}$; $t_{2}$: $a_{2}$.
Task:
  1. (1)
    Build a unified set of UNIQUE answers by merging semantically identical strings across $a_{0}$, $a_{1}$, $a_{2}$.
  2. (2)
    Compute a GLOBAL frequency vector over the unique answers based on how many of $a_{0}$$a_{1}$$a_{2}$ map to each canonical answer.
  3. (3)
    ASSIGN probabilities for EACH interventions ($t_{0}$, $t_{1}$, $t_{2}$) to be EXACTLY this global frequency vector normalised by 3 (counts/3), after duplicate aggregation and before rounding.
Re-scoring Structure:
Merging Rules
  • Ignore case, whitespace, punctuation, trivial formatting, and plural/singular differences.
  • Choose a clean canonical form for answers (e.g., title case).
  • If multiple inputs ($a_{0}$$a_{1}$$a_{2}$) map to the same canonical answer, aggregate them by SUMMING that answer’s frequency before normalisation.
Probability rules (HARD CONSTRAINTS)
  • Let count[i] be how many of $a_{0}$$a_{1}$$a_{2}$ map to answers[i]. Then for every intervention T in $t_{0}$, $t_{1}$, $t_{2}$, set T[i] = round(count[i]/3, 2).
  • Do NOT output equal probabilities across answers when counts differ.
  • Each probability $\in[0.00,1.00]$; sums may be ¡ 1.00 after rounding.
  • If an intervention has no plausible answers, use zeros. Return EXACTLY one line of STRICT JSON, no wrapper, no extra text.
@@ Appendix B Case Study
Case Study I on WebQSP
Question: what is the capital of argentina?
Ground-truth: Buenos Aires
Step 1: Counterfactual Prompting
  • initial prompt $t_{0}$:
@@
Step 2: Panel-based Re-scoring
You are given 3 candidate answers for the same question from different prompts: $t_{0}$: $a_{0}$; $t_{1}$: $a_{1}$; $t_{2}$: $a_{2}$.
Task: (1) Build a unified set of UNIQUE answers by merging semantically identical strings across $a_{0}$, $a_{1}$, $a_{2}$. (2) Compute a GLOBAL frequency vector over the unique answers based on how many of $a_{0}$$a_{1}$$a_{2}$ map to each canonical answer. (3) Assign probabilities for each scenario ($t_{0}$, $t_{1}$, $t_{2}$) to be exactly this global frequency vector normalised by 3 (counts/3), after duplicate aggregation and before rounding. Merging rules: ignore case, whitespace, punctuation, and plural/singular differences; choose a clean canonical form for answers; if multiple inputs map to the same canonical answer, aggregate by summing before normalisation. Probability rules: A) Let count[i] be how many of $a_{0}$, $a_{1}$, $a_{2}$} map to answers[i]. Then for every scenario S in $t_{0}$, $t_{1}$, $t_{2}$}, set S[i] = round(count[i]/3, 2). B) Do not output equal probabilities across answers when counts differ. C) Each probability $\in[0.00,1.00]$; sums may be <1.00<1.00 after rounding. D) If a scenario has no plausible answers, use zeros. Return exactly one line of strict JSON, no wrapper, no extra text.
Canonical merge: Buenos Aires, Argentina
@@
"Buenos Aires": 0.62, "Argentina": 0.38
Step 3: Accurate Answer Generation
Panel JSON:
@@
Causal effect variation ($\text{CE}_{\text{var}}$):
$\text{CE}_{\text{var}}$ $(\text{Buenos Aires})$
$=\max(0.62,0.68,0.80)-\min(0.62,0.68,0.80)$
$=0.80-0.62=0.18$
$\text{CE}_{\text{var}}$ $(\text{Argentina})$
$=\max(0.38,0.32,0.20)-\min(0.38,0.32,0.20)$
$=0.38-0.20=0.18$
Average causal effect ($\overline{\text{CE}}$):
$\overline{\text{CE}}(\text{Buenos Aires})=(0.62+0.68+0.80)/3=0.70$
$\overline{\text{CE}}(\text{Argentina})=(0.38+0.32+0.20)/3=0.30$
Causal calibration index (CCI):
$\text{CCI}(\text{Buenos Aires})=0.70\times(1-0.18)=0.70\times 0.82=0.57$
$\text{CCI}(\text{Argentina})=0.30\times(1-0.18)=0.30\times 0.82=0.25$
Final Answer: {Buenos Aires}
Case Study II on MetaQA
Question: The films that share directors with the films [Tiresia] are written by who?
Ground-truth: Bertrand Bonello
Step 1: Counterfactual Prompting
  • initial prompt $t_{0}$:
@@
Step 2: Panel-based Re-scoring
You are given 3 candidate answers for the same question from different prompts: $t_{0}$: $a_{0}$; $t_{1}$: $a_{1}$; $t_{2}$: $a_{2}$.
Task: (1) Build a unified set of UNIQUE answers by merging semantically identical strings across $a_{0}$, $a_{1}$, $a_{2}$. (2) Compute a GLOBAL frequency vector over the unique answers based on how many of $a_{0}$$a_{1}$$a_{2}$ map to each canonical answer. (3) Assign probabilities for each scenario ($t_{0}$, $t_{1}$, $t_{2}$) to be exactly this global frequency vector normalised by 3 (counts/3), after duplicate aggregation and before rounding. Merging rules: ignore case, whitespace, punctuation, and plural/singular differences; choose a clean canonical form for answers; if multiple inputs map to the same canonical answer, aggregate by summing before normalisation. Probability rules: A) Let count[i] be how many of $a_{0}$, $a_{1}$, $a_{2}$} map to answers[i]. Then for every scenario S in $t_{0}$, $t_{1}$, $t_{2}$}, set S[i] = round(count[i]/3, 2). B) Do not output equal probabilities across answers when counts differ. C) Each probability $\in[0.00,1.00]$; sums may be <1.00<1.00 after rounding. D) If a scenario has no plausible answers, use zeros. Return exactly one line of strict JSON, no wrapper, no extra text.
Canonical merge: Bertrand Bonello, Luca Fazzi
@@
"Bertrand Bonello": 0.67, "Luca Fazzi": 0.33
Step 3: Accurate Answer Generation
Panel JSON:
@@
Causal effect variation ($\text{CE}_{\text{var}}$):
$\text{CE}_{\text{var}}$ $(\text{Bertrand Bonello})$
$=\max(0.67,0.67,0.67)-\min(0.67,0.67,0.67)$
$=0.00$
$\text{CE}_{\text{var}}$ $(\text{Luca Fazzi})$
$=\max(0.33,0.33,0.33)-\min(0.33,0.33,0.33)$
$=0.00$
Average causal effect ($\overline{\text{CE}}$):
$\overline{\text{CE}}(\text{Bertrand Bonello})$ $=(0.67+0.67+0.67)/3=0.67$
$\overline{\text{CE}}(\text{Luca Fazzi})$ $=(0.33+0.33+0.33)/3=0.33$
Causal calibration index (CCI):
$\text{CCI}(\text{Bertrand Bonello})$ $=0.67\times(1-0.00)=0.67$
$\text{CCI}(\text{Luca Fazzi})$ $=0.33\times(1-0.00)=0.33$
Final Answer: {Bertrand Bonello}